OpenStudy (anonymous):

chuck travels 215 miles dana travels 200 miles if chuck rate is 3mph more than dana what speed is chuck traveling?

OpenStudy (anonymous):

help me

OpenStudy (nowhereman):

question is, do they travel those distances in the same time?

OpenStudy (dumbcow):

d=r*t chuck: 215 = (x+3)t 215/(x+3) = t dana: 200 = xt 200/x = t set them equal to each other and solve for x 215/(x+3) = 200/x cross multiply 215x = 200(x+3) 215x = 200x +600 15x = 600 x = 40

OpenStudy (anonymous):

i had 43 thank you mr.cow i refuse to call you dumb

OpenStudy (dumbcow):

haha thank you, 43 would be correct for chucks speed, 40 is danas speed